# Yongtao Li, Lihua Feng, Yuejian Peng

### v1

@@ 2.1 A spectral Lovász–Simonovits theorem
When we embed a graph $H$ with $q$ edges into a vertex part of the bipartite Turán graph $T_{n,2}$, we can compute that the spectral radius of the resulting graph will be decreasing in the order of star, complete graph, complete bipartite graph, cycle, path and matching. In addition, we can verify that if we embed two matchings of size $q_{1}$ and $q_{2}$ with $q_{1}+q_{2}=q$ into two vertex parts, respectively, then the resulting graph has larger spectral radius than that of $Y_{n,2,q}$. On the contrary, if we embed two stars of sizes $q_{1}$ and $q_{2}$ into two parts, respectively, then the resulting graph has spectral radius smaller than that of $T_{n,2,q}$.
2.2 A spectral extension on a stability result
In 2021, Xiao and Katona [56] initiated the study on stability of Theorem 1.1. The triangle covering number $\tau_{3}(G)$ is defined as the minimum number of vertices in a set that contains at least one vertex of each triangle in $G$. In particular, $\tau_{3}(G)=1$ if and only if all triangles of $G$ share a common vertex. It was shown in [56] that if $G$ has more than $\lfloor n^{2}/4\rfloor$ edges and $\tau_{3}(G)\geq 2$, then $G$ has at least $n-2$ triangles. Furthermore, they proposed a problem to study the stability of Theorem 1.2. Recently, this problem was solved by Liu and Mubayi [32], as well as Balogh and Clemen [3] independently.
Before stating their result, we need to define two extremal graphs. Firstly, let $G_{1}$ be the graph of order $n$, where $V(G_{1})=A\cup B$ with $|A|=\lceil{n}/{2}\rceil+a$ and $|B|=\lfloor{n}/{2}\rfloor-a$, where $a\in\mathbb{N}$ is chosen later. Pick $2(s-1)$ vertices $x_{1},y_{1},\ldots,x_{s-1},y_{s-1}$ in $A$ and two vertices $u_{1},u_{2}$ in $B$. Adding $s$ edges $\{u_{1},u_{2}\}$ and $\{x_{i},y_{i}\}$ into $K_{|A|,|B|}$, and delete $\alpha$ edges $\{u_{1},x_{1}\},\ldots,\{u_{1},x_{\alpha}\}$, where $\alpha:=s-t-a^{2}-\mathbf{1}_{\{2\nmid n\}}a$. Then $e(G_{1})=\lfloor n^{2}/4\rfloor+t$ and $\tau_{3}(G_{1})=s$. Secondly, let $G_{2}$ be the graph with the same vertex partition. Unlike the previous construction, we now pick $2s$ vertices $x_{1},y_{1},\ldots,x_{s},y_{s}\in A$ and one vertex $u\in B$, add the edges $\{x_{i},y_{i}\}$ into $A$, and delete the edges $\{u,x_{1}\},\ldots,\{u,x_{\alpha}\}$ between $A$ and $B$. Similarly, we have $e(G_{2})=\lfloor n^{2}/4\rfloor+t$ and $\tau_{3}(G_{2})=s$. It is easy to see that $t(G_{1})=(s-1)|B|+|A|-2\alpha$ and $t(G_{2})=s|B|-\alpha$. Now, we choose an appropriate $a\in\mathbb{N}$ to minimize the number of triangles in the above constructions.
@@ 2.2 A spectral extension on a stability result
An independent work was also studied by Liu and Mubayi [32, Theorem 1.4]. In other words, there are two ways to embed $s$ disjoint edges into the partition in order to minimize the number of triangles, i.e., embedding $s-1$ or $s$ edges into the larger vertex part, and other edges into the smaller part. Recently, Li, Feng and Peng [27] presented a spectral version of the result of Xiao and Katona by showing that if $G$ is an $n$-vertex graph with $\lambda(G)\geq\lambda(T_{n,2})$ and $\tau_{3}(G)\geq 2$, then $G$ contains at least $n-3$ triangles. This bound is the best possible and it provides a spectral stability of Ning–Zhai’s result in Theorem 1.3.
Next, we shall consider the general case $\tau_{3}(G)\geq s$ for every $s\in\mathbb{N}$. The third main result in this paper reveals an asymptotically spectral extension on Theorem 2.4.
@@
If $G$ is a graph on $n$ vertices with $\lambda(G)\geq\lambda(T_{n,2})$ and $\tau_{3}(G)\geq s$, then for every $n\geq 113s^{2}$, there are at least $s{n}/{2}-5s^{2}$ triangles in $G$.
@@
Remark. Theorem 2.5 gives a stability of Theorem 2.3. By applying the fact $\lambda(G)\geq 2e(G)/n$, we can see that $e(G)>e(T_{n,2})$ implies $\lambda(G)>\lambda(T_{n,2})$. As a by-product of Theorem 2.5, we obtain that if $e(G)>e(T_{n,2})$ and $\tau_{3}(G)\geq s$, then $G$ has $s{n}/{2}-O(1)$ triangles. This result could be viewed as an asymptotic version of Theorem 2.4, since $t(G_{1})\approx t(G_{2})\approx s{n}/{2}-O(1)$. With additional efforts, one can improve the constant in Theorem 2.5, but it seems difficult to determine the exact bound or the spectral extremal graphs corresponding to the edge-extremal graphs $G_{1}$ and $G_{2}$ in the general case.
@@
Our approach. The main idea in our proofs of Theorems 2.2, 2.3 and 2.5 uses a very simple yet surprisingly powerful technique known as the supersaturation-stability; see Theorem 3.3. Informally, every graph with few triangles can be made bipartite by deleting a quite small number of edges. The word “quite small” means a small constant number of edges, say $6q$ in Lemma 4.1, which is independent of the order $n$. This is slightly different from the classical stability method, which requires us to delete roughly $\varepsilon n^{2}$ edges with a small number $\varepsilon>0$. Although this might sound somewhat vague, we will make it more precise in Section 4. Unlike the traditional stability method, it is sufficient to utilize the condition $n\geq 300q^{2}$ only, instead of requiring that $q$ is fixed and $n$ is sufficiently large (or a tower-type function on $q$), since our proof avoids the use of the celebrated triangle removal lemma***The lemma says that for each $\varepsilon>0$, there is $\delta>0$ such that every graph on $n$ vertices with at most $\delta n^{3}$ triangles can be made triangle-free by removing at most $\varepsilon n^{2}$ edges. That is, a graph with a subcubic number of triangles can be made triangle-free by removing a subquadratic number of edges; see [13].. This allows us to put $q$ in a wide range that can be growth with $n$. Hence, we substantially improve the lower bound on the order $n$ of a graph with respect to $q$. We believe that the supersaturation-stability method used in this paper may be interesting in its own right, and perhaps could have other applications elsewhere, just as the classical stability approach has now seen widespread use. Last but not least, apart from using the supersaturation-stability, we shall make use of the spectral technique due to Cioabă, Feng, Tait and Zhang [9] as well as the double-eigenvector technique applied by Fang, Tait and Zhai [19].
@@
Organization. In Section 3, we shall introduce some preliminaries, including the spectral triangle counting lemma and the supersaturation-stability; see Lemma 3.2 and Theorem 3.3, respectively. Intuitively it is clear that if a graph $G$ has large spectral radius and it is far from being bipartite, then it is unlikely that all the triangles of $G$ are located on few edges of $G$. To warm up, we shall give a “strong” stability result for graphs with few triangles and large spectral radius; see Theorem 3.4. In Section 4, we will outline the sketch of our method and then we present the detailed proof of Theorem 2.2. In Section 5, we shall use the double-eigenvector technique to give the proof of Theorem 2.3. The proof of Theorem 2.5 will be provided in Section 6. At the end of this paper, we will propose some related spectral extremal graph problems for interested readers.
@@
Notation. We usually write $G=(V,E)$ for a graph with vertex set $V=\{v_{1},\ldots,v_{n}\}$ and edge set $E=\{e_{1},\ldots,e_{m}\}$, where each $e_{i}$ is a two-element subset of $V$. Here, we admit $|V|=n$ and $|E|=m$. If $S\subseteq V$ is a subset of the vertex set, then $G[S]$ denotes the subgraph of $G$ induced by $S$, i.e., the graph on $S$ whose edges are those edges of $G$ with both endpoints in $S$. For abbreviation, we denote $e(S)=e(G[S])$. We will write $G[S,T]$ for the induced subgraph of $G$ whose edges have one endpoint in $S$ and the other in $T$, and similarly, we write $e(S,T)$ for the number of edges of $G[S,T]$. For every $s,t\in\mathbb{N}$, we write $K_{s,t}$ for the complete bipartite graph with parts of sizes $s$ and $t$. Let $N(v)$ be the set of vertices adjacent to a vertex $v$ and let $d(v)=|N(v)|$. Moreover, we denote $N_{S}(v)=N(v)\cap S$ and $d_{S}(v)=|N_{S}(v)|$ for simplicity. We will write $\lambda(G)$ for the spectral radius of $G$ and $t(G)$ for the number of triangles in $G$. Sometimes, we write $\lambda$ and $t$ for short. The triangle covering number $\tau_{3}(G)$ is defined as the minimum number of vertices that hit all triangles of $G$. For a vertex set $S\subseteq V(G)$, we write $G\setminus S$ for the subgraph of $G$ by deleting all vertices of $S$ and their incident edges. Let $\mathbf{x}=(x_{1},x_{2},\ldots,x_{n})^{\top}$ be a Perron vector of $G$, i.e., a non-negative eigenvector corresponding to $\lambda(G)$. By scaling, we may assume further that $\max\{x_{v}:v\in V(G)\}=1$. For a subset $S\subseteq V(G)$, we denote $x_{S}=\sum_{v\in S}x_{v}$ for notational convenience. In particular, we have $x_{V}=\sum_{v\in V(G)}x_{v}$.
@@ 3 Preliminaries
Note that in the case $q=1$, we have $Y_{n,2,1}=T_{n,2,1}$.
Lemma 3.1 (See [25]).
(a) If $n$ is even, then $\lambda(Y_{n,2,1})$ is the largest root of
$f(x)=x^{3}-x^{2}-(n^{2}x)/4+{n^{2}}/{4}-n.$
(b) If $n$ is odd, then $\lambda(Y_{n,2,1})$ is the largest root of
$g(x)=x^{3}-x^{2}+{(1-n^{2})x}/{4}+{n^{2}}/{4}-n+{3}/{4}.$
Consequently, we have
@@
$\lambda^{2}(Y_{n,2,1})>\left\lfloor{n^{2}}/{4}\right\rfloor+2.$

### v2

@@ 2.1 A spectral Lovász–Simonovits theorem
Remark. When embedding a graph $H$ with $q$ edges into one partite set of $T_{n,2}$, we can verify that the spectral radius of the resulting graph decreases according to the following order of $H$: star, complete graph, complete bipartite graph (that is not a star), cycle, path, and matching. Moreover, it can be verified that if we embed two matchings of sizes $q_{1}$ and $q_{2}$ (with $q_{1}+q_{2}=q$) into two partite sets, respectively, then the resulting graph has larger spectral radius than $Y_{n,2,q}$. In contrast, if we embed two stars of sizes $q_{1}$ and $q_{2}$ into two partite sets, respectively, then the resulting graph has smaller spectral radius than $T_{n,2,q}$.
2.2 A spectral extension of a stability result
The triangle covering number $\tau_{3}(G)$ is the minimum number of vertices in a set that contains at least one vertex of each triangle in $G$. In particular, $\tau_{3}(G)=1$ if and only if all triangles of $G$ share a common vertex. Xiao and Katona [59] gave a stability of Theorem 1.1 by showing that if $e(G)>\lfloor n^{2}/4\rfloor$ and $\tau_{3}(G)\geq 2$, then $G$ has at least n2n-2 triangles. Moreover, they proposed a problem to study the stability of Theorem 1.2. Recently, this problem was solved by Liu and Mubayi [34], as well as Balogh and Clemen [3] independently.
Before stating their result, we need to define two extremal graphs. Firstly, let $G_{1}$ be the graph of order $n$, where $V(G_{1})=A\cup B$ with $|A|=\lceil{n}/{2}\rceil+a$ and $|B|=\lfloor{n}/{2}\rfloor-a$, where $a\in\mathbb{N}$ is chosen later. Pick 2(s1)2(s-1) vertices $x_{1},y_{1},\ldots,x_{s-1},y_{s-1}$ in $A$ and two vertices $u_{1},u_{2}$ in $B$. Adding $s$ edges $\{u_{1},u_{2}\}$ and $\{x_{i},y_{i}\}$, for $i\in\{1,\ldots,s-1\}$ into $K_{|A|,|B|}$, and delete $\alpha$ edges $\{u_{1},x_{1}\},\ldots,\{u_{1},x_{\alpha}\}$, where $\alpha:=s-t-a^{2}-\mathbf{1}_{\{2\nmid n\}}a$. Then $e(G_{1})=\lfloor n^{2}/4\rfloor+t$ and $\tau_{3}(G_{1})=s$. Secondly, let $G_{2}$ be the graph with the same vertex partition. Unlike the previous construction, we now pick $2s$ vertices $x_{1},y_{1},\ldots,x_{s},y_{s}\in A$ and one vertex $u\in B$, add the $s$ edges $\{x_{i},y_{i}\}$, for $i\in\{1,\ldots,s\}$ into $A$, and delete the $\alpha$ edges $\{u,x_{1}\},\ldots,\{u,x_{\alpha}\}$ between $A$ and $B$. Similarly, we have $e(G_{2})=\lfloor n^{2}/4\rfloor+t$ and $\tau_{3}(G_{2})=s$. It is easy to see that $t(G_{1})=(s-1)|B|+|A|-2\alpha$ and $t(G_{2})=s|B|-\alpha$. Now, we choose an appropriate $a\in\mathbb{N}$ to minimize the number of triangles in the above constructions.
@@ 2.2 A spectral extension of a stability result
An independent work was also studied by Liu and Mubayi [34, Theorem 1.4]. In other words, there are two ways to embed $s$ disjoint edges in the partition in order to minimize the number of triangles, that is, embedding s1s-1 or $s$ edges in the larger vertex part, and other edges in the smaller part. Recently, Li, Feng and Peng [29] presented a spectral version of the result of Xiao and Katona [59] by showing that if $G$ is an $n$-vertex graph with $\lambda(G)\geq\lambda(T_{n,2})$ and $\tau_{3}(G)\geq 2$, then $G$ contains at least n3n-3 triangles. This bound is the best possible and provides a spectral stability of the Ning–Zhai result in Theorem 1.3.
Next, we shall consider the general case $\tau_{3}(G)\geq s$ for every $s\in\mathbb{N}$. The third main result in this paper reveals an asymptotically spectral extension of Theorem 2.4.
@@
If $G$ is a graph with $n\geq 28s^{2}$ vertices with $\lambda(G)\geq\lambda(T_{n,2})$ and $\tau_{3}(G)\geq s$, then there are at least $\frac{1}{2}sn-5s^{2}$ triangles in $G$.
@@
Remark. Theorem 2.5 gives the stability of Theorem 2.3. Applying the fact $\lambda(G)\geq\frac{2e(G)}{n}$, we see that $e(G)>e(T_{n,2})$ implies $\lambda(G)>\lambda(T_{n,2})$. As a by-product of Theorem 2.5, we find that if $e(G)>e(T_{n,2})$ and $\tau_{3}(G)\geq s$, then $G$ contains $\frac{1}{2}sn-O(1)$ triangles. This result could be viewed as an asymptotic version of Theorem 2.4, since $t(G_{1})\approx t(G_{2})\approx\frac{1}{2}sn-O(1)$. With additional efforts, the constant term of Theorem 2.5 can be slightly improved, but it seems difficult to establish the exact bound and determine the spectral extremal graphs corresponding to the extremal graphs $G_{1}$ and $G_{2}$ in Theorem 2.4.
2.3 Our approach and organization
@@
Our approach. The main idea in our proofs of Theorems 2.2, 2.3 and 2.5 uses a simple yet surprisingly powerful technique known as the supersaturation-stability method. Informally, every graph with few triangles can be made bipartite by deleting a relatively small number of edges. Here, “relatively small” means a constant number of edges independent of $n$; see Theorem 3.6. This is different from the classical stability method, which typically requires the removal of roughly $\varepsilon n^{2}$ edges for some small $\varepsilon>0$. Although this description may sound somewhat vague, we will make it more precise in Section 4. Unlike the traditional stability method, our method requires only the condition $n\geq 121q^{2}$, instead of requiring that $q$ is fixed and $n$ is sufficiently large (or a tower-type function in $q$). This is because our proof avoids the use of the celebrated triangle removal lemma***The lemma says that for each $\varepsilon>0$, there is $\delta>0$ such that every graph on $n$ vertices with at most $\delta n^{3}$ triangles can be made triangle-free by removing at most $\varepsilon n^{2}$ edges. That is, a graph with a subcubic number of triangles can be made triangle-free by removing a subquadratic number of edges; see [13].. Consequently, $q$ is allowed to vary over a wide range and may even grow with $n$. Hence, we significantly improve the lower bound on the order $n$ of a graph with respect to $q$. We believe that the supersaturation-stability method used in this paper may be of independent interest and could potentially find further applications elsewhere, much like the classical stability approach, which has now seen widespread use. Last but not least, in addition to the supersaturation-stability method, we also employ the spectral technique developed by Cioabă, Feng, Tait and Zhang [9], as well as the double-eigenvector technique applied by Fang, Tait and Zhai [19].
@@
Organization. In Section 3, we show the tightness of $q$ in Theorem 2.2 and introduce preliminaries for counting triangles. As a warm up, we give a spectral stability result for graphs with few triangles; see Theorem 3.6. In Section 4, we outline the main ideas of our method and then present the proof of Theorem 2.2. In Section 5, we employ the double-eigenvector technique to prove Theorem 2.3. The proof of Theorem 2.5 is provided in Section 6. Finally, we propose several related spectral extremal problems for readers.
@@
Notation. We write $G=(V,E)$ for a graph with vertex set $V=\{v_{1},\ldots,v_{n}\}$ and edge set $E=\{e_{1},\ldots,e_{m}\}$, where we admit $|V|=n$ and $|E|=m$. If $S\subseteq V$ is a subset of the vertex set, then $G[S]$ denotes the subgraph of $G$ induced by $S$, i.e., the graph on $S$ whose edges are those edges of $G$ with both endpoints in $S$. For abbreviation, we denote $e(S)=e(G[S])$. We write $G[S,T]$ for the induced subgraph of $G$ whose edges have one endpoint in $S$ and the other in $T$, and we write $e(S,T)=e(G[S,T])$. We write $K_{s,t}$ for the complete bipartite graph with parts of sizes $s$ and $t$. Let $N(v)$ be the set of vertices adjacent to a vertex $v$ and let $d(v)=|N(v)|$. For simplicity, we denote $N_{S}(v)=N(v)\cap S$ and $d_{S}(v)=|N_{S}(v)|$. We write $\lambda(G)$ for the spectral radius of $G$ and $t(G)$ for the number of triangles in $G$. Sometimes, we write $\lambda$ and $t$ for short. The triangle covering number $\tau_{3}(G)$ is defined as the minimum number of vertices that hit all triangles of $G$. For a vertex set $S\subseteq V(G)$, we write $G\setminus S$ for the subgraph of $G$ by deleting all vertices of $S$ and their incident edges. Let $\bm{x}=(x_{1},x_{2},\ldots,x_{n})^{\mathrm{T}}$ be a Perron vector of $G$, i.e., a non-negative eigenvector corresponding to $\lambda(G)$. By scaling, we may further assume that $\max\{x_{v}:v\in V(G)\}=1$. For a subset $S\subseteq V(G)$, we denote $x_{S}=\sum_{v\in S}x_{v}$. In particular, we have $x_{V}=\sum_{v\in V(G)}x_{v}$.
@@ 3 Preliminaries
Recall that $Y_{n,2,q}$ is a graph obtained from the bipartite Turán graph $T_{n,2}$ by embedding $q$ pairwise disjoint edges into the partite set of size $\lceil n/2\rceil$. In what follows, we give an estimate on the spectral radius of $Y_{n,2,q}$.
Lemma 3.1.
(a) If $n$ is even, then $\lambda(Y_{n,2,q})$ is the largest root of
$f_{1}(x)=x^{3}-x^{2}-(n^{2}x)/4+n^{2}/4-qn.$
(b) If $n$ is odd, then $\lambda(Y_{n,2,q})$ is the largest root of
$f_{2}(x)=x^{3}-x^{2}-(n^{2}x)/4+x/4+n^{2}/4-1/4+q-qn.$
@@
Consequently, we have $\sqrt{\left\lfloor{n^{2}}/{4}\right\rfloor+2q}<\lambda(Y_{n,2,q})<\frac{n}{2}+\frac{2q}{n}+\frac{4q}{n^{2}}$ for every $n\geq 4q$.
Proof.
(a) Let $\bm{x}=(x_{1},x_{2},\ldots,x_{n})^{\mathrm{T}}$ be the Perron eigenvector of $\lambda(Y_{n,2,q})$. Assume that $V(Y_{n,2,q})=X_{1}\cup X_{2}\cup Y$, where $X_{1}=\{u_{1},v_{1},u_{2},v_{2},\ldots,u_{q},v_{q}\}$ is the set of vertices of the matching of size $q$, $X_{1}\cup X_{2}$ and $Y$ are partite sets of $K_{\frac{n}{2},\frac{n}{2}}$. Since two vertices with the same neighborhood have the same value in corresponding coordinates of $\bm{x}$, we may assume that $x_{u_{i}}=x_{v_{i}}=a$ for $1\leq i\leq q$, $x_{u}=b$ for each $u\in X_{2}$, and $x_{w}=c$ for each $w\in Y$. Then
$\begin{cases}\lambda a=a+\frac{n}{2}c,\\ \lambda b=\frac{n}{2}c,\\ \lambda c=2qa+(\frac{n}{2}-2q)b.\end{cases}$
Thus, $\lambda(Y_{n,2,q})$ is the largest eigenvalue of
$B_{1}=\begin{bmatrix}1&0&\frac{n}{2}\\ 0&0&\frac{n}{2}\\ 2q&\frac{n}{2}-2q&0\end{bmatrix}$.
By calculation, we know that $\lambda(Y_{n,2,q})$ is the largest root of
$f_{1}(x):=\det(xI-B_{1})=x^{3}-x^{2}-(n^{2}x)/4+n^{2}/4-qn.$
(b) Let $V(Y_{n,2,q})=X_{1}\cup X_{2}\cup Y$ be a partition, where $X_{1}=\{u_{1},v_{1},\ldots,u_{q},v_{q}\}$ spans a matching with $q$ edges, $X_{1}\cup X_{2}$ and $Y$ are partite sets of $K_{\frac{n+1}{2},\frac{n-1}{2}}$ satisfying $|X_{1}|+|X_{2}|=\frac{n+1}{2}$ and $|Y|=\frac{n-1}{2}$. A similar argument yields that $\lambda(Y_{n,2,q})$ is the largest eigenvalue of
$B_{2}=\begin{bmatrix}1&0&\frac{n-1}{2}\\ 0&0&\frac{n-1}{2}\\ 2q&\frac{n+1}{2}-2q&0\end{bmatrix}$.
Thus, $\lambda(Y_{n,2,q})$ is the largest root of
$f_{2}(x):=\det(xI-B_{2})=x^{3}-x^{2}+x/4-(n^{2}x)/4+n^{2}/4-1/4-qn+q.$
Moreover, it is easy to check that
$f_{1}(\sqrt{n^{2}/4+2q})=2q\sqrt{n^{2}/4+2q}-qn-2q<0$
and
$f_{2}(\sqrt{(n^{2}-1)/4+4})=2q\sqrt{(n^{2}-1)/4+2q}-qn-q<0.$
Consequently, we get $\sqrt{\lfloor n^{2}/4\rfloor+4}<\lambda(Y_{n,2,q})$. Similarly, we can verify that $f_{1}(\frac{n}{2}+\frac{2q}{n}+\frac{4q}{n^{2}})>0$ and $f_{2}(\frac{n}{2}+\frac{2q}{n}+\frac{4q}{n^{2}})>0$, which implies that $\lambda(Y_{n,2,q})<\frac{n}{2}+\frac{2q}{n}+\frac{4q}{n^{2}}$, as desired. ∎
Similar to the proof of Lemma 3.1, we prove the following proposition.
Proposition 3.2.
For any $q\geq 2$, we have $\lambda(Y_{n,2,q})<\lambda(T_{n,2,q})$.
Proof.
We start the proof in two cases. (a) If $n$ is even, by a simple computation as in the proof of Lemma 3.1, we can obtain that $\lambda(T_{n,2,q})$ is the largest root of
$h_{1}(x)$ $:=x^{4}-x^{3}-(n^{2}x^{2})/4-qx^{2}+(n^{2}x)/4-(nx)/2-nqx$
$\quad-(nq)/2+(n^{2}q)/4-(nq^{2})/2.$
Let $f_{1}(x)$ be the polynomial defined in Lemma 3.1. It follows that
$xf_{1}(x)-h_{1}(x)=qx^{2}+(nx)/2+(nq^{2})/2+(nq)/2-(n^{2}q)/4.$
It is easy to see that for every $x>\frac{n}{2}$, we have
$xf_{1}(x)-h_{1}(x)>\tfrac{n}{2}f_{1}(\tfrac{n}{2})-h_{1}(\tfrac{n}{2})=\tfrac{1}{4}(n^{2}+2q(1+q)n)>0.$
We conclude that $xf_{1}(x)>h_{1}(x)$ for every $x>\frac{n}{2}$. Note that $\lambda(Y_{n,2,q})$ is the largest root of $xf_{1}(x)$. Combining with $\lambda(Y_{n,2,q}),\lambda(T_{n,2,q})>\frac{n}{2}$, we obtain $\lambda(Y_{n,2,q})<\lambda(T_{n,2,q})$.
(b) If $n$ is odd, then $\lambda(T_{n,2,q})$ is the largest root of
$h_{2}(x)$ $:=x^{4}-x^{3}+x^{2}/4-(n^{2}x^{2})/4-qx^{2}+x/4-(nx)/2+(n^{2}x)/4+qx-nqx$
$\quad+q/4-(nq)/2+(n^{2}q)/4+q^{2}/2-(nq^{2})/2.$
Let $f_{2}(x)$ be the polynomial defined in Lemma 3.1. Then
$xf_{2}(x)-h_{2}(x)=qx^{2}+(n-1)x/2+(nq)/2-(n^{2}q)/4+(n-1)q^{2}/2-q/4.$
For every $x>\frac{n-1}{2}$, it follows that
$xf_{2}(x)-h_{2}(x)>\tfrac{n-1}{2}f_{2}(\tfrac{n-1}{2})-h_{2}(\tfrac{n-1}{2})=\tfrac{1}{4}(n-1)(2q^{2}+n-1)>0.$
In this case, observe that $\lambda(Y_{n,2,q}),\lambda(T_{n,2,q})>\frac{n-1}{2}$, so we get $\lambda(Y_{n,2,q})<\lambda(T_{n,2,q})$. ∎
3.1 The tightness of $q$ in Theorem 2.2
In this section, we provide an example showing that the bound of $q$ in Theorem 2.2 is tight up to a constant factor. Recall that $Y_{n,2,q}$ and $T_{n,2,q}$ are obtained from $T_{n,2}$ by adding a matching, and a star with $q$ edges into the partite set of size $\lceil n/2\rceil$, respectively.
Example 3.3.
If $q=0.8\sqrt{n}$ and $n\geq 110$ are integers, then $\lambda(T_{n,2,q-1})>\lambda(Y_{n,2,q})$. However, the graph $T_{n,2,q-1}$ contains exactly $(q-1)\lfloor n/2\rfloor$ triangles.
Proof.
We partition the proof into two cases. (a) If $n$ is even, by a similar argument as in the proof of Lemma 3.1, we know that $\lambda(T_{n,2,q-1})$ is the largest root of
$g_{1}(x)$ $:=x^{4}-x^{3}-(n^{2}x^{2})/4-qx^{2}+x^{2}+(n^{2}x)/4+(nx)/2-nqx$
$\quad-(n^{2}/4)+(nq)/2+(n^{2}q)/4-(nq^{2})/2.$
Let $f_{1}(x)$ be defined in Lemma 3.1. Since $q=0.8\sqrt{n}$, by computation, we see that
$xf_{1}(x)-g_{1}(x)=(0.8\sqrt{n}-1)x^{2}-0.5nx-0.2n^{5/2}+0.57n^{2}-0.4n^{3/2}.$
For every $n\geq 4$ and x>n/2x>{n}/{2}, it follows that
$\frac{\mathrm{d}}{\mathrm{d}x}(xf_{1}(x)-g_{1}(x))=(1.6\sqrt{n}-2)x-0.5n>0.$
Then for every $x>n/2$, we have
$xf_{1}(x)-g_{1}(x)\geq\tfrac{n}{2}f_{1}(\tfrac{n}{2})-g_{1}(\tfrac{n}{2})=0.07n^{2}-0.4n^{3/2}>0.$
We conclude that $xf_{1}(x)>g_{1}(x)$ for every $x>n/2$ and $n\geq 33$. Note that $\lambda(Y_{n,2,q})$ and $\lambda(T_{n,2,q-1})$ are the largest roots of $xf_{1}(x)$ and $g_{1}(x)$, respectively. Since $\lambda(Y_{n,2,q})>\frac{n}{2}$ and $\lambda(T_{n,2,q-1})>\frac{n}{2}$, we obtain $\lambda(Y_{n,2,q})<\lambda(T_{n,2,q-1})$.
(b) If $n$ is odd, by computation, we get that $\lambda(T_{n,2,q-1})$ is the largest root of
$g_{2}(x)$ $=x^{4}-x^{3}+(5x^{2})/4-(n^{2}x^{2})/4-qx^{2}-(3x)/4+(nx)/2+(n^{2}x)/4+qx-nqx$
$\quad+1/4-n^{2}/4-(3q)/4+(nq)/2+(n^{2}q)/4+q^{2}/2-(nq^{2})/2.$
Let $f_{2}(x)$ be determined in Lemma 3.1. Substituting $q=0.8\sqrt{n}$, we compute that
[…]
For every $x>\frac{n-1}{2}$, we have
$\frac{\mathrm{d}}{\mathrm{d}x}(xf_{2}(x)-g_{2}(x))=(1.6\sqrt{n}-2)x-0.5(n-1)>0.$
It follows that for every $n\geq 110$ and $x>\frac{n-1}{2}$,
$xf_{2}(x)-g_{2}(x)>\tfrac{n-1}{2}f_{2}(\tfrac{n-1}{2})-g_{2}(\tfrac{n-1}{2})=0.07n^{2}-0.8n^{3/2}+0.68n+0.8\sqrt{n}-0.75>0.$
Thus, we get $xf_{2}(x)>g_{2}(x)$ for every $x>\frac{n-1}{2}$. It is easy to see that $\lambda(Y_{n,2,q})>\frac{n-1}{2}$ and $\lambda(T_{n,2,q-1})>\frac{n-1}{2}$. So we conclude that $\lambda(Y_{n,2,q})<\lambda(T_{n,2,q-1})$. ∎
3.2 Counting the number of triangles